# Curious crossing-critical edges – variations on an example of S̆irán̆

### v1

@@ Abstract.
Motivated by Kuratowski’s theorem, a Kuratowski subgraph of a graph is a subgraph that is a subdivided $K_{5}$ or a subdivided $K_{3,3}$. An edge is crossing-critical if the crossing number decreases after removing the edge. In this note, we present the following examples: a graph with an edge that is crossed in every optimal drawing of the graph, but the edge is not in any Kuratowski subgraph of the graph; a graph with an edge that is in every Kuratowski subgraph but is not crossed in any optimal drawing of the graph; and a graph with a crossing-critical edge that is not present in any Kuratowski subgraph and is not crossed in any optimal drawing of the graph.
@@ 1. Introduction
A graph is planar if it can be drawn in the plane without crossing edges, i.e., such that the images of two edges only have their common endvertices in common; in other words edges do not cross in the drawing. The celebrated Kuratowski theorem [2] states that a graph if planar precisely when it does not contain a subdivision of the complete graph $K_{5}$ or the complete bipartite graph $K_{3,3}$. $H$ is called a Kuratowski subgraph of the graph $G$ if $H$ is a subgraph of $G$ and $H$ is a subdivision of either $K_{3,3}$ or $K_{5}$. Note that Kuratowski subgraphs are simple graphs by definition. In a sense, Kuratowski theorem states that Kuratowski subgraphs are forcing crossings to appear in any drawing of the graph.
The crossing number $\operatorname{cr}(\mathcal{D})$ of a drawing $\mathcal{D}$ of $G$ is $\sum\limits_{e,f\in E(G)}|\mathcal{D}(e)^{0}\cap\mathcal{D}(f)^{0}|$, where for an edge $e$, $\mathcal{D}(e)^{0}$ is the interior of $\mathcal{D}(e)$ , and the summation goes for unordered pairs of distinct edges. The crossing number $\operatorname{cr}(G)$ of a graph $G$ is the minimum crossing number over all its drawings. A drawing $\mathcal{D}$ of $G$ is called optimal if $\operatorname{cr}(\mathcal{D})=\operatorname{cr}(G)$. An edge $e\in E(G)$ is called crossing-critical if $\operatorname{cr}(G-e)<\operatorname{cr}(G)$.
@@
If $G$ is a non-planar graph and $h$ is an edge that belongs to every Kuratowski subgraph of $G$, then $h$ is crossing-critical, as by Kuratowski’s theorem $\operatorname{cr}(G-h)=0$.
@@
    If $G$ is a non-planar graph with an edge $h$ such that $h$ is in every Kuratowski subgraph, must $G$ have an optimal drawing in which $h$ is crossed?
@@
The answer to all questions in Problem 1 is in the negative, i.e.,
@@
    A graph can have an edge that is crossed in every optimal drawing, but is not in any Kuratowski subgraph of the graph. (An example is the graph is $G_{1}$ and the unique $s$-$t$ edge $e$.)
@@
    A graph can have an edge that is in every Kuratowski subgraph but is not crossed in any optimal drawing of the graph. (An example is the graph is $G_{2}$ and the unique $v_{1}$-$v_{3}$ edge $f$.)
@@
Furthermore, all three negative answers can be provided with simple graphs.
We note that the “furthermore” part of the previous claim is obvious: in each of the $G_{i}$, subdivide all but one parallel copies of each $x$-$y$ edge to obtain a simple graph. The edge $e$ does not have parallel copies, so it is never subdivided, and the edge $f$ does not have parallel copies in $G_{2}$. Hence these simple graphs also have the required properties.
@@ 2. Proof of our claims
Let $H$ be a Kuratowski subgraph of $G_{0}$. This $H$ is a subdivided $K_{3,3}$. If $H$ has no subdividing vertices, then there is a vertex $y\in V(G_{0})$ such that $H$ is a spanning subgraph of $G_{0}-y$. As $H$ is $3$-regular, $y$ is not adjacent to any degree $3$ vertex in $G_{0}$. Therefore $y\in\{v_{2},v_{4}\}$. As for $j\in\{2,4\}$ the graph $G_{0}-v_{j}$ is $3$-regular, we now have $H=G-v_{j}$. However, $G-v_{j}$ contains the $3$-cycle $uv_{1}v_{3}$ and therefore it is not $K_{3,3}$. Thus, $H$ must have a unique subdividing vertex $w$.
@@
Figure 3. The drawing $\mathcal{D}_{4}$ of $G_{0}$.

### v2

@@ Abstract.
Motivated by Kuratowski’s theorem, a Kuratowski subgraph of a graph is a subgraph that is a subdivided $K_{5}$ or a subdivided $K_{3,3}$. An edge is crossing-critical if the crossing number decreases after removing the edge. In this note, we present the following examples: a graph with an edge that is crossed in every optimal drawing of the graph, but the edge is not in any Kuratowski subgraph of the graph; a graph with an edge that is in every Kuratowski subgraph but is not crossed in any optimal drawing of the graph; and a graph with a crossing-critical edge that is not present in any Kuratowski subgraph and is not crossed in any optimal drawing of the graph. Fáry’s theorem implies that the Kuratowski subgraphs are the only obstructions to a graph having a crossing-free drawing with all edges drawn as straight lines. The three example graphs given also hold if we restrict drawings to only have straight line edges, and thus also apply to the rectilinear crossing number.
@@ 1. Introduction
A graph is planar if it can be drawn in the plane without crossing edges, i.e., such that the images of two edges only have their common endvertices in common; in other words edges do not cross in the drawing. The celebrated Kuratowski theorem [3] states that a graph if planar precisely when it does not contain a subdivision of the complete graph $K_{5}$ or the complete bipartite graph $K_{3,3}$. $H$ is called a Kuratowski subgraph of the graph $G$ if $H$ is a subgraph of $G$ and $H$ is a subdivision of either $K_{3,3}$ or $K_{5}$. Note that Kuratowski subgraphs are simple graphs by definition. In a sense, Kuratowski theorem states that Kuratowski subgraphs are forcing crossings to appear in any drawing of the graph. Fáry’s theorem (proved independently in [2, 6]) states that every planar graph can be drawn without crossings even if we restrict to drawings using straight line edges (the rectilinear drawings). Thus, the Kuratowski subgraphs are also the only obstacles to having planar straight line drawings.
The crossing number $\operatorname{cr}(\mathcal{D})$ of a drawing $\mathcal{D}$ of $G$ is $\sum\limits_{e,f\in E(G)}|\mathcal{D}(e)^{0}\cap\mathcal{D}(f)^{0}|$, where for an edge $e$, $\mathcal{D}(e)^{0}$ is the interior of $\mathcal{D}(e)$ , and the summation goes for unordered pairs of distinct edges. The crossing number $\operatorname{cr}(G)$ of a graph $G$ is the minimum crossing number over all its drawings. A drawing $\mathcal{D}$ of $G$ is called optimal if $\operatorname{cr}(\mathcal{D})=\operatorname{cr}(G)$. An edge $e\in E(G)$ is called crossing-critical if $\operatorname{cr}(G-e)<\operatorname{cr}(G)$. The rectilinear crossing number $\overline{\operatorname{cr}}(G)$ of the graph $G$ is the minimum crossing number of its rectilinear drawings. Clearly $\operatorname{cr}(G)\leq\overline{\operatorname{cr}}(G)$, with equality precisely when $G$ has an optimal drawing that is rectilinear. An edge $e\in E(G)$ called crossing-critical with respect to the rectilinear crossing number if $\overline{\operatorname{cr}}(G-e)<\overline{\operatorname{cr}}(G)$.
@@
If $G$ is a non-planar graph and $h$ is an edge that belongs to every Kuratowski subgraph of $G$, then $h$ is crossing-critical, as by Kuratowski’s theorem $\operatorname{cr}(G-h)=0$, and by Fáry’s theorem this edge is also crossing critical with respect to the rectilinear crossing number.
@@
    If $G$ is a graph with an edge $h$ such that $h$ is in every Kuratowski subgraph, must $G$ have an optimal drawing in which $h$ is crossed?
@@
The answer to all questions in Problem 1 is in the negative (see Figure 2 for the examples), i.e.,
@@
    A graph can have an edge that is crossed in every optimal drawing, but is not in any Kuratowski subgraph of the graph. (An example is the graph $G_{1}$ and the unique $s$-$t$ edge $e$.)
@@
    A non-planar graph can have an edge that is in every Kuratowski subgraph but is not crossed in any optimal drawing of the graph. (An example is the graph $G_{2}$ and the unique $v_{1}$-$v_{3}$ edge $f$.)
@@
Furthermore, all three negative answers can be provided with simple graphs. Moreover, the these simple graphs provide the corresponding examples for the rectilinear crossing number as well.
TWe note that the “furthermore” part of the previous claim is obvious: in each of the $G_{i}$, subdivide all but one of the parallel copies of each $x$-$y$ edge to obtain a simple graph $G_{i}^{*}$. The edge $e$ does not have parallel copies, so it is never subdivided, and the edge $f$ does not have parallel copies in $G_{2}$. Hence these simple graphs also have the required properties. The “moreover” part follows from the fact that we have straight line drawings of the underlying simple graphs $G_{i}$ (Figure 3) used in the proof, and the simple graphs $G_{i}^{*}$ have corresponding straight line drawing: draw the not subdivided copy of a multi-edge as a straight line, and by placeing the subdivideding copies of this edges of the multi-edges close to the original edge.
@@ 2. Proof of our claims
Let $H$ be a Kuratowski subgraph of $G_{0}$. This $H$ is a subdivided $K_{3,3}$. First, suppose $H$ has no subdividing vertices. Then there is a vertex $y\in V(G_{0})$ such that $H$ is a spanning subgraph of $G_{0}-y$. As $H$ is $3$-regular, $y$ is not adjacent to any degree $3$ vertex in $G_{0}$. Therefore $y\in\{v_{2},v_{4}\}$. As for $j\in\{2,4\}$ the graph $G_{0}-v_{j}$ is $3$-regular, so we now have $H=G-v_{j}$. However, $G-v_{j}$ contains the $3$-cycle $uv_{1}v_{3}$ and therefore it is not $K_{3,3}$. Thus, $H$ must have a unique subdividing vertex $w$.
@@
[…]
Figure 3. The rectilinear drawing $\mathcal{D}_{4}$ of $G_{0}$, and a rectilinear drawing equivalent with $\mathcal{D}_{0}$.